• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Holmstrom

    Log entry time 00:02:34 on August11,2005

    Entry number 151422

    keyword=Swing Shift Summary

    16:00 Shift started with MCC doing beam studies in the Hall.
    
    19:00 Talked to MCC they said that beam would be back in more than 1 hour. 
    
    
    21:00 Talked to MCC "they are working on it".
    
    
    21:20 MCC says that we will get low CW beam.
    
    21:49 Run flipper IWHP IN
    
    21:50 Beam returns, positions not good they are adjusting. 
    
    22:23 Positions look good ask for production currents.
    
    Production Runs  IWHP IN
    4278  22:24-23:21  100k
    4279  23:40-
    
    23:21 MCC asks us to move the tager to empty, so they can see the beam on the 
    dump.
    
    23:33 MCC sees the beam on the dump. Moving target back to He.
    
    23:54 Septum IOC is lost.
    
    24:00 Shift ends rebooting IOCs seems to be coming back.